OpenStudy (twizttiez):

Solve, using the substitution method. y – 3x = 7 21 + 9x = 3y A. There is no solution. B. There are an infinite number of solutions. C. The solution is (1, 10) D. The solution is (28, 0)

OpenStudy (danjs):

solve the first equation for y y = 3x + 7

OpenStudy (danjs):

place that y into the second equation

OpenStudy (danjs):

21 + 9x = 3*(3x + 7)

OpenStudy (twizttiez):

I think B

OpenStudy (danjs):

i didnt solve it , idk

OpenStudy (anonymous):

So, from where he left off, distribute the 3out

OpenStudy (anonymous):

You'll get 21+9x=9x+21

OpenStudy (twizttiez):

y – 3x = 7 y – 3x+3x = 7+3x y + 0x = 3x + 7 y + 0 = 3x + 7 y = 3x + 7 Therefore infinite number of solutions

OpenStudy (anonymous):

That's the exact same line isn't it?

OpenStudy (anonymous):

Good job. It's the exact same line, therefore there are an infinite number.

OpenStudy (twizttiez):

YAY!

OpenStudy (danjs):

It would be infinite solutions because the second equations is a constant scalar multiple of the first (-3)

OpenStudy (twizttiez):

Could you help with one more?

OpenStudy (anonymous):

Sure!